# Radiative decays of $P$-wave bottom baryons from light-cone sum rules

### v1

@@ III Decays of flavor $\mathbf{\bar{3}}_{F}$ baryons
Following the same procedures, we proceed to investigate the four $SU(3)$ flavor $\mathbf{\bar{3}}_{F}$ bottom baryon multiplets, $[\mathbf{\bar{3}}_{F},1,0,\lambda]$, $[\mathbf{\bar{3}}_{F},0,1,\rho]$, $[\mathbf{\bar{3}}_{F},1,1,\rho]$, and $[\mathbf{\bar{3}}_{F},2,1,\rho]$, in the following subsections.
III.2 The $[\mathbf{\bar{3}}_{F},1,0,\lambda]$ doublet
The $[\mathbf{\bar{3}}_{F},1,0,\lambda]$ doublet contains $\Lambda_{b}^{0}({1/2}^{-})$, $\Lambda_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{0}({1/2}^{-})$, $\Xi_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{-}({1/2}^{-})$, and $\Xi_{b}^{-}({3/2}^{-})$. We study their radiative transitions into ground-state bottom baryons, and derive the following non-zero coupling constants:
$g_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $1.90^{+0.86}_{-0.39}~{}\rm GeV^{-1}\,,$ (20)
$g_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $1.10^{+0.53}_{-0.22}~{}\rm GeV^{-1}\,,$
$g_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $1.40^{+0.51}_{-0.38}~{}\rm GeV^{-1}\,,$
$g_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.55^{+0.20}_{-0.13}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $1.70^{+0.58}_{-0.47}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.97^{+0.27}_{-0.24}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $1.20^{+0.41}_{-0.30}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.47^{+0.17}_{-0.11}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.095^{+0.061}_{-0.070}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.055^{+0.037}_{-0.040}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.067^{+0.049}_{-0.049}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.026^{+0.017}_{-0.018}~{}\rm GeV^{-1}\,.$
We can further derive the following radiative decay widths:
$\Gamma_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $340^{+6200}_{-~{}340}~{}\rm keV\,,$ (21)
$\Gamma_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $110^{+3400}_{-~{}~{}83}~{}\rm keV\,,$
$\Gamma_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $210^{+3500}_{-~{}210}~{}\rm keV\,,$
$\Gamma_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $9.2^{+250.0}_{-~{}~{}8.1}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $1200^{+11000}_{-~{}1200}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $550^{+6400}_{-~{}550}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $660^{+5800}_{-~{}660}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $38^{+430}_{-~{}38}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $3.8^{+36.2}_{-~{}3.8}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $1.7^{+20.9}_{-~{}1.7}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $2.1^{+20.6}_{-~{}2.1}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.12^{+1.34}_{-0.12}~{}\rm keV\,.$
As examples, we plot in Fig. 3 some of the coupling constants with respect to the Borel mass $T$. The results of the $\Lambda_{b}$ and $\Xi_{b}$ baryons are separately summarized in Table 2 and Table 3, together with their total strong decay widths calculated in Ref. Tan:2023opd .
Figure 3: Coupling constants with respect to the Borel mass T: (a) $g_{\Lambda^{0}_{b}[\frac{1}{2}^{-}]\to\Sigma_{b}^{0}\gamma}$, (b) $g_{\Xi^{0}_{b}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime 0}\gamma}$, (c) $g_{\Xi^{-}_{b}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime-}\gamma}$, (d) $g_{\Lambda^{0}_{b}[\frac{3}{2}^{-}]\to\Sigma_{b}^{0}\gamma}$, (e) $g_{\Xi^{0}_{b}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime 0}\gamma}$, and (f) $g_{\Xi^{-}_{b}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime-}\gamma}$. Here the bottom baryon doublet $[\mathbf{\bar{3}}_{F},1,0,\lambda]$ is investigated.
III.3 The $[\mathbf{\bar{3}}_{F},0,1,\rho]$ singlet
The $[\mathbf{\bar{3}}_{F},0,1,\rho]$ singlet contains $\Lambda_{b}^{0}({1/2}^{-})$, $\Xi_{b}^{0}({1/2}^{-})$, and $\Xi_{b}^{-}({1/2}^{-})$. We study their radiative transitions into ground-state bottom baryons, and derive the following non-zero coupling constants:
$g_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $1.20^{+0.57}_{-0.49}~{}\rm GeV^{-1}\,,$ (22)
$g_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $1.50^{+0.64}_{-0.69}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $1.00^{+0.38}_{-0.30}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $1.10^{+0.41}_{-0.31}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.10^{+0.01}_{-0.02}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.12^{+0.01}_{-0.03}~{}\rm GeV^{-1}\,.$
We can further derive the following radiative decay widths:
$\Gamma_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $170^{+2700}_{-~{}130}~{}\rm keV\,,$ (23)
$\Gamma_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $280^{+6800}_{-~{}280}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $440^{+390}_{-440}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $720^{+640}_{-720}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $4.4^{+1.5}_{-4.0}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $8.5^{+6.9}_{-8.4}~{}\rm keV\,.$
As examples, we plot in Fig. 4 some of the coupling constants with respect to the Borel mass $T$. The results of the $\Lambda_{b}$ and $\Xi_{b}$ baryons are separately summarized in Table 2 and Table 3, together with their total strong decay widths calculated in Ref. Tan:2023opd .
Figure 4: Coupling constants with respect to the Borel mass T: (a) $g_{\Lambda^{0}_{b}[\frac{1}{2}^{-}]\to\Sigma_{b}^{0}\gamma}$, (b) $g_{\Xi^{0}_{b}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime 0}\gamma}$, (c) $g_{\Xi^{-}_{b}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime-}\gamma}$, (d) $g_{\Lambda^{0}_{b}[\frac{1}{2}^{-}]\to\Sigma_{b}^{*0}\gamma}$, (e) $g_{\Xi^{0}_{b}[\frac{1}{2}^{-}]\to\Xi_{b}^{*0}\gamma}$, and (f) $g_{\Xi^{-}_{b}[\frac{1}{2}^{-}]\to\Xi_{b}^{*-}\gamma}$. Here the bottom baryon singlet $[\mathbf{\bar{3}}_{F},0,1,\rho]$ is investigated.
Figure 5: Coupling constants with respect to the Borel mass T: (a) $g_{\Lambda^{0}_{b}[\frac{1}{2}^{-}]\to\Lambda_{b}^{0}\gamma}$, (b) $g_{\Xi^{0}_{b}[\frac{1}{2}^{-}]\to\Xi_{b}^{0}\gamma}$, (c) $g_{\Xi^{-}_{b}[\frac{1}{2}^{-}]\to\Xi_{b}^{-}\gamma}$, (d) $g_{\Lambda^{0}_{b}[\frac{3}{2}^{-}]\to\Lambda_{b}^{0}\gamma}$, (e) $g_{\Xi^{0}_{b}[\frac{3}{2}^{-}]\to\Xi_{b}^{0}\gamma}$, and (f) $g_{\Xi^{-}_{b}[\frac{3}{2}^{-}]\to\Xi_{b}^{-}\gamma}$. Here the bottom baryon doublet $[\mathbf{\bar{3}}_{F},1,1,\rho]$ is investigated.
Figure 6: Coupling constants with respect to the Borel mass T: (a) $g_{\Lambda^{0}_{b}[\frac{3}{2}^{-}]\to\Sigma_{b}^{0}\gamma}$, (b) $g_{\Xi^{0}_{b}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime 0}\gamma}$, (c) $g_{\Xi^{-}_{b}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime-}\gamma}$, (d) $g_{\Lambda^{0}_{b}[\frac{5}{2}^{-}]\to\Sigma_{b}^{*0}\gamma}$, (e) $g_{\Xi^{0}_{b}[\frac{5}{2}^{-}]\to\Xi_{b}^{*0}\gamma}$, and (f) $g_{\Xi^{-}_{b}[\frac{5}{2}^{-}]\to\Xi_{b}^{*-}\gamma}$. Here the bottom baryon doublet $[\mathbf{\bar{3}}_{F},2,1,\rho]$ is investigated.
III.4 The $[\mathbf{\bar{3}}_{F},1,1,\rho]$ doublet
The $[\mathbf{\bar{3}}_{F},1,1,\rho]$ doublet contains $\Lambda_{b}^{0}({1/2}^{-})$, $\Lambda_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{0}({1/2}^{-})$, $\Xi_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{-}({1/2}^{-})$, and $\Xi_{b}^{-}({3/2}^{-})$. We study their radiative transitions into ground-state bottom baryons, and derive the following non-zero coupling constants:
$g_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Lambda_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.12^{+0.02}_{-0.04}~{}\rm GeV^{-1}\,,$ (24)
$g_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.10^{+0.04}_{-0.04}~{}\rm GeV^{-1}\,,$
$g_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.057^{+0.019}_{-0.021}~{}\rm GeV^{-1}\,,$
$g_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Lambda_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.16^{+0.06}_{-0.05}~{}\rm GeV^{-1}\,,$
$g_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.070^{+0.023}_{-0.027}~{}\rm GeV^{-1}\,,$
$g_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.080^{+0.029}_{-0.029}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.13^{+0.04}_{-0.06}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.10^{+0.03}_{-0.05}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.052^{+0.020}_{-0.021}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.18^{+0.06}_{-0.07}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.064^{+0.026}_{-0.026}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.074^{+0.030}_{-0.031}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{-}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.20^{+0.07}_{-0.05}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.004^{+0.003}_{-0.004}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.005^{+0.002}_{-0.002}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{-}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.29^{+0.10}_{-0.10}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.006^{+0.002}_{-0.002}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.007^{+0.002}_{-0.002}~{}\rm GeV^{-1}\,.$
We can further derive the following radiative decay widths:
$\Gamma_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Lambda_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $34^{+64}_{-30}~{}\rm keV\,,$ (25)
$\Gamma_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.10^{+11.10}_{-~{}0.10}~{}\rm keV\,,$
$\Gamma_{\Lambda_{b}^{0}[\frac{1}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.32^{+2.95}_{-0.32}~{}\rm keV\,,$
$\Gamma_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Lambda_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $65^{+130}_{-~{}55}~{}\rm keV\,,$
$\Gamma_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.60^{+2.60}_{-0.60}~{}\rm keV\,,$
$\Gamma_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.24^{+3.26}_{-0.24}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $41^{+78}_{-41}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $3.5^{+17.9}_{-~{}3.5}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{1}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $1.2^{+8.1}_{-1.2}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $85^{+160}_{-~{}85}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $1.7^{+7.9}_{-1.7}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.84^{+4.89}_{-0.84}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{-}[\frac{1}{2}^{+}]\gamma}$ $=$ $97^{+200}_{-~{}91}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.01^{+0.03}_{-0.01}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{1}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.01^{+0.08}_{-0.01}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{-}[\frac{1}{2}^{+}]\gamma}$ $=$ $220^{+420}_{-210}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $1.5^{+6.8}_{-1.5}~{}\rm keV\,,$
$\Gamma_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.75^{+5.19}_{-0.75}~{}\rm keV\,.$
As examples, we plot in Fig. 5 some of the coupling constants with respect to the Borel mass $T$. T The results of the $\Lambda_{b}$ and $\Xi_{b}$ baryons are separately summarized in Table 2 and Table 3, together with their total strong decay widths calculated in Ref. Tan:2023opd .
III.5 The $[\mathbf{\bar{3}}_{F},2,1,\rho]$ doublet
The $[\mathbf{\bar{3}}_{F},2,1,\rho]$ doublet contains $\Lambda_{b}^{0}({3/2}^{-})$, $\Lambda_{b}^{0}({5/2}^{-})$, $\Xi_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{0}({5/2}^{-})$, $\Xi_{b}^{-}({3/2}^{-})$, and $\Xi_{b}^{-}({5/2}^{-})$. We study their radiative transitions into ground-state bottom baryons, and derive the following non-zero coupling constants:
$g_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.69^{+0.31}_{-0.25}~{}\rm GeV^{-1}\,,$ (26)
$g_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.28^{+0.14}_{-0.11}~{}\rm GeV^{-1}\,,$
$g_{\Lambda_{b}^{0}[\frac{5}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.23^{+0.11}_{-0.10}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime 0}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.57^{+0.26}_{-0.20}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{3}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.22^{+0.13}_{-0.10}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{0}[\frac{5}{2}^{-}]\to\Xi_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.18^{+0.10}_{-0.10}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{\prime-}[\frac{1}{2}^{+}]\gamma}$ $=$ $0.048^{+0.014}_{-0.012}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{3}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.021^{+0.010}_{-0.010}~{}\rm GeV^{-1}\,,$
$g_{\Xi_{b}^{-}[\frac{5}{2}^{-}]\to\Xi_{b}^{*-}[\frac{3}{2}^{+}]\gamma}$ $=$ $0.014^{+0.003}_{-0.003}~{}\rm GeV^{-1}\,.$
We can further derive the following radiative decay widths:
$\Gamma_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{0}[\frac{1}{2}^{+}]\gamma}$ $=$ $73^{+520}_{-~{}73}~{}\rm keV\,,$ (27)
$\Gamma_{\Lambda_{b}^{0}[\frac{3}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $3.8^{+38.2}_{-~{}3.8}~{}\rm keV\,,$
$\Gamma_{\Lambda_{b}^{0}[\frac{5}{2}^{-}]\to\Sigma_{b}^{*0}[\frac{3}{2}^{+}]\gamma}$ $=$ $11^{+84}_{-11}~{}\rm keV\,,$
[…]

### v3

@@ III Decays of flavor $\mathbf{\bar{3}}_{F}$ baryons
Following the same procedures, we proceed to investigate the four $SU(3)$ flavor $\mathbf{\bar{3}}_{F}$ bottom baryon multiplets, $[\mathbf{\bar{3}}_{F},1,0,\lambda]$, $[\mathbf{\bar{3}}_{F},0,1,\rho]$, $[\mathbf{\bar{3}}_{F},1,1,\rho]$, and $[\mathbf{\bar{3}}_{F},2,1,\rho]$:
  • The $[\mathbf{\bar{3}}_{F},1,0,\lambda]$ doublet contains $\Lambda_{b}^{0}({1/2}^{-})$, $\Lambda_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{0}({1/2}^{-})$, $\Xi_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{-}({1/2}^{-})$, and $\Xi_{b}^{-}({3/2}^{-})$.
  • The $[\mathbf{\bar{3}}_{F},0,1,\rho]$ singlet contains $\Lambda_{b}^{0}({1/2}^{-})$, $\Xi_{b}^{0}({1/2}^{-})$, and $\Xi_{b}^{-}({1/2}^{-})$.
  • The $[\mathbf{\bar{3}}_{F},1,1,\rho]$ doublet contains $\Lambda_{b}^{0}({1/2}^{-})$, $\Lambda_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{0}({1/2}^{-})$, $\Xi_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{-}({1/2}^{-})$, and $\Xi_{b}^{-}({3/2}^{-})$.
  • The $[\mathbf{\bar{3}}_{F},2,1,\rho]$ doublet contains $\Lambda_{b}^{0}({3/2}^{-})$, $\Lambda_{b}^{0}({5/2}^{-})$, $\Xi_{b}^{0}({3/2}^{-})$, $\Xi_{b}^{0}({5/2}^{-})$, $\Xi_{b}^{-}({3/2}^{-})$, and $\Xi_{b}^{-}({5/2}^{-})$.
We study their radiative transitions into ground-state bottom baryons, and the obtained sum rule equations are listed in the supplementary file “OPE.nb”. The results of the $\Lambda_{b}$ and $\Xi_{b}$ baryons are separately summarized in Table 2 and Table 3, respectively.
Table 2: Radiative decay widths of the $P$-wave bottom baryons $\Lambda_{b}$ belonging to the $SU(3)$ flavor $\mathbf{\bar{3}}_{F}$ representation, calculated using the light-cone sum rule method within the framework of heavy quark effective theory. The total radiative decay widths (T.R.W.) are listed in the seventh column.
Multiplets Baryon       Mass     Difference    Decay channels   Coupling constants   Decay width     T.R.W.
($j^{P}$) (GeV) (MeV) ($\rm GeV^{-1}$) (keV) (keV)
  $[\mathbf{\bar{3}}_{F},0,1,\rho]$   $\Lambda_{b}({1\over 2}^{-})$ $5.92^{+0.17}_{-0.19}$ $\Lambda_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $1.20^{+0.57}_{-0.49}$ $170^{+2700}_{-~{}130}$ $450^{+9500}_{-~{}410}$
$\Lambda_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $1.50^{+0.64}_{-0.69}$ $280^{+6800}_{-~{}280}$
$[\mathbf{\bar{3}}_{F},1,1,\rho]$ $\Lambda_{b}({1\over 2}^{-})$ $5.92^{+0.13}_{-0.10}$ $7\pm 3$ $\Lambda_{b}^{0}({1\over 2}^{-})\to\Lambda_{b}^{0}\gamma$ $0.12^{+0.02}_{-0.04}$ $34^{+64}_{-30}$ $34^{+78}_{-30}$
$\Lambda_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $0.10^{+0.04}_{-0.04}$ $0.10^{+11.10}_{-~{}0.10}$
$\Lambda_{b}^{0}({1\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.057^{+0.019}_{-0.021}$ $0.32^{+2.95}_{-0.32}$
$\Lambda_{b}({3\over 2}^{-})$ $5.92^{+0.13}_{-0.10}$ $\Lambda_{b}^{0}({3\over 2}^{-})\to\Lambda_{b}^{0}\gamma$ $0.16^{+0.06}_{-0.05}$ $65^{+130}_{-~{}55}$ $65^{+140}_{-~{}56}$
$\Lambda_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $0.070^{+0.023}_{-0.027}$ $0.60^{+2.60}_{-0.60}$
$\Lambda_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.080^{+0.029}_{-0.029}$ $0.24^{+3.26}_{-0.24}$
$[\mathbf{\bar{3}}_{F},2,1,\rho]$ $\Lambda_{b}({3\over 2}^{-})$ $5.93^{+0.13}_{-0.13}$ $17\pm 7$ $\Lambda_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{0}\gamma$ $0.69^{+0.31}_{-0.25}$ $73^{+520}_{-~{}73}$ $77^{+560}_{-~{}77}$
$\Lambda_{b}^{0}({3\over 2}^{-})\to\Sigma_{b}^{*0}\gamma$ $0.28^{+0.14}_{-0.11}$ $3.8^{+38.2}_{-~{}3.8}$
[…]